OpenStudy (anonymous):

How would i factor f(x)=x^3-3x^2-16x-12 given that f=6 f(6)=0 ?

OpenStudy (anonymous):

\[f(6)=x ^{3}-3x ^{2}-16x-12\]

OpenStudy (anonymous):

Because f(6)=0, you know that 6 is a solution of the function. THerefore, you get (x-6)(something)=x^3-3x^2-16x-12=0

OpenStudy (anonymous):

YOu can then divide the big polynomial by (x-6) and then solve for the rest of the solutions

OpenStudy (anonymous):

Wouldn't I divide x^3-3x^2-16x-12=0 by (x-6) ?

OpenStudy (anonymous):

that's what I meant by "the big polynomial"

OpenStudy (anonymous):

Oh, ok. Thank you!

OpenStudy (anonymous):

Do I use synthetic division or the regular division?

OpenStudy (anonymous):

doesn't matter. Whichever is easier

OpenStudy (anonymous):

Ok, Thanks

OpenStudy (anonymous):

Regular

OpenStudy (anonymous):

they both do the same thing